The risk-free rate is 5 percent. Stock A has a beta 1.2 and Stock B has a beta 1.4. Stock A has a required return of 11 percent. What is Stock B’s required return?
1
Expert's answer
2020-05-07T18:48:51-0400

Stock required return formula is:

"rs = rf + b\u00d7(rm - rf)," where rf is risk-free rate, b is beta, rm is market rate.

For stock A:

"0.11 = 0.05 + 1.2\u00d7(rm - 0.05),"

rm = 0.1.

Stock B’s required return is:

"rs = 0.05 + 1.4\u00d7(0.1 - 0.05) = 0.12"or 12 percent.